[Brahms-l] Shift report 20040103 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sat Jan 03 2004 - 15:55:17 EST
------------------------------------------------------------
Date: Sat Jan  3 15:55:16 2004, Shift: 08:00-16:00
Supervisor: Oana Ristea, with: 
Log book#15 pages 22 - 22

Shift Summary:
------------------------------------------------------------
1. No beam for entire shift.

2. Run 9337 - pedestal run for detectors;
   run 9338 - TPC pedestal run.

3. Flemming fixed the TDC missing channels for TOFW (see e-log).
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 3 15:55:38 2004

This archive was generated by hypermail 2.1.8 : Sat Jan 03 2004 - 15:55:53 EST